OpenStudy (anonymous):

Solve the open sentence. –2 ≤ n + 4 ≤ 7 A. n ≥ 6 and n ≤ –4 B. n ≤ 2 and n ≤ 7 C. n ≤ –6 and n ≤ 4 D. n ≥ –6 and n ≤ 3

OpenStudy (anonymous):

\[ –2 ≤ n + 4 ≤ 7\] subtract \(4\) all the way across get \[-6\leq n\leq 3\]

OpenStudy (anonymous):

what the monkey is an "open sentence"????

OpenStudy (anonymous):

subtract 4 we get - 2 -4 <=n+4-4 <=7-4 -6<= n<=3 that is n =>-6 and n<=3

OpenStudy (anonymous):

btw why they write one as a compound inequality and the other as \[n ≥ –6 \text{ and } n ≤ 3\] is anyones guess but that is the right answer

OpenStudy (anonymous):

A. n ≥ 6 and n ≤ –4 B. n ≤ 2 and n ≤ 7 C. n ≤ –6 and n ≤ 4 D. n ≥ –6 and n ≤ 3 those are the choices so these are open sentences I guess

OpenStudy (anonymous):

thanks @satellite73